OpenStudy (anonymous):

Quadrilateral ABCD is a rhombus with diagonals that intersect at point E. Given angle ABE = 30 degrees and segment AD = 5, what is the area of triangle EBC?

OpenStudy (anonymous):

 it is a rhombus so ab=bc=cd=ad=5 and from the graph ABE=CBE=30degrees so \[BE=5*\sin 30=5\sqrt{3}/2\] \[CE=5*\cos 30=5/2\] area ebc=\[0.5*(5\sqrt{3}/2)*(5/2)=25\sqrt{3}/8\]